OpenStudy (anonymous):

(7 minus 2 square root of 5)(5 plus 4 square root of 5) HELLPPPPPP

OpenStudy (anonymous):

35+18sqrt5-40 =-5+18sqrt5

OpenStudy (campbell_st):

\[(7 - 2\sqrt{5})(5+4\sqrt{5})\] its binomial expansion lots of ways to do it  collect like terms